# thanks: Part of the Instituto Avanzado de Cosmología (IAC) Collaboration

### v1

@@ II.5 Exact solution
See Figure 2 for a 3D representation of the spherical ground state, the first excited state, and the sum of these two states of the SFDM. The main result of this work is that we obtained, in an analytical way, a spherical ground state, as expected, and a first excited state with a two-lobes shape towards the north and south of the center of the configuration. This structure is like that of the hydrogen atom, called a gravitational atom. The two-lobes structure can explain why satellite galaxies in MW-like galaxies are not distributed spherically, but rather along a north-south line, which has been called VPOS. This behavior has not only been observed in the Milky Way (Pawlowski and Kroupa, 2013, 2020), but also in Andromeda (Ibata et al., 2013; Conn et al., 2013) and Centaurus A (Müller et al., 2018, 2021). In this way, our result has the potential to explain this anisotropic distribution of satellite galaxies. Additionally, we tested the model with the distribution of satellites in 6 other galaxies reported in Nashimoto et al. (2022).
@@
Figure 2: (Top) A numerical simulation of the ground state of the SFDM. (Center) The same simulation but showing the evolution of the two-lobes excited state. (Bottom) The two states combined in the same image.
Now, from (27) and (28), the physical densities are given by $\rho_{k}(r)=R_{k}(r)R_{k}^{*}(r)$, and we define $\rho_{0}:=R_{00}^{2}$ and $\rho_{1}:=R_{10}^{2}$, with $\rho_{0}$ the central density of the ground state, and $\rho_{1}$ the corresponding maximum density of the first excited state (see Figure 3):
$\rho_{0}(r)=\rho_{0}\frac{\sin^{2}x}{x^{2}},$ (29)
$\rho_{1}(r)=\rho_{1}\left(\frac{\sin x}{x^{2}}-\frac{\cos x}{x}\right)^{2}.$ (30)
Notice that the density of the ground state (29) is of the same form as in the case of the real SF solution (equation (4)), for $j=1$ and $l=\pi/R$. In our case, the value $R=\pi/l$, define the radius of the first minima of the ground state (see Figure 3). In (Robles and Matos, 2013a), the authors defined $R$ as the radius of the SFDM halo. However, in this work, since there is still contribution to the mass beyond the first minimum, as seen in Figure 3, we define the maximum radius of every SF state, $R_{i}^{\text{max}}$, such that
$\rho_{i}(R_{i}^{\text{max}})=200\rho_{c},$ (31)
@@
Another difference with the previous work (Robles and Matos, 2013a), is that the SFDM states in the case of the real SF are all centered at r=0, only their wavelength decreases and their amplitude changes, being all spherically symmetric. In that case, all the excited states share the definition of the radius $R$ of the halo. In the present work, for the complex SF, the second excited state has a different functional form, which gives it the two-lobes shape, with $\rho_{1}(0)=0$ and its maximum density outside $r=0$ (see Figure 3).
@@
$M_{0}(r)=\frac{2\pi\rho_{0}}{l^{3}}\left(x-\cos x\sin x\right),$ (32)
$M_{1}(r)=\frac{2\pi\rho_{1}}{l^{3}}\left(\frac{\cos(2x)}{x}-\frac{1}{x}+\cos x\sin x+x\right).$ (33)
@@
Figure 3: The purple dashed line is the density of the ground state, centered at $r=0$. The blue solid line is the density of the first excited state, with its first maximum outside $r=0$. Note that the ground state decays very quickly, and the first excited state decays slower, even if its first maximum is smaller than the first maximum of the ground state.
III Rotation curves of galaxies
@@
To test this model and obtain the best values of the parameters $\rho_{0}$, $\rho_{1}$ and $l$ for the ground and the first excited states of the SFDM halo, respectively, we fit the rotation curves of the Milky Way, Andromeda, Centaurus A, as the main galaxies where the anisotropy of the satellite galaxies has been observed. Also, we fit the rotation curves of 6 MW-like galaxies whose satellite galaxies have been observed (Nashimoto et al., 2022).
@@
The rotation curve (the tangential velocity of the satellites) is simply $v^{2}=GM(r)/r$, for $M$ the mass of the galaxy. In what follows, we fit the rotation curves of different galaxies with this formula, such that
$v(r)=\sqrt{\frac{G}{r}\left(M_{\text{b}}+M_{\text{d}}+M_{0}+M_{1}\right)},$ (34)
@@
where the subscripts are for the bulge, disc, and ground and first excited states, respectively. We do not consider the gas contribution.
@@
To test the plausibility of the model in explaining the anisotropic distribution of satellite galaxies around a host galaxy, we fit the best models that reproduce the observed rotation curve, with the ground state (spherically symmetric) and the first excited state (with two lobes) of the mSFDM as the DM halo. As said before, we define the radius of every state such that $\rho_{i}(R_{i})=200\rho_{c}$, for $i=0,1$, where $\rho_{c}=1.283\times 10^{-7}M_{\odot}\text{pc}^{-3}$ is the critical density of the Universe. Both states present density oscillations until reaching the density of $200\rho_{c}$. These halos are like concentric shells that contribute less and less to the gravitational potential.
@@
For the Milky Way, Andromeda, and Centaurus A, we fit the baryonic components and the mSFDM halo. We assume that the motion of stars and gas in the galaxies is circular, without considering non-axisymmetric structures such as the bar and spiral arms. For the MW, we fit an exponential bulge, an exponential disc, and the DM halo, with 7 free parameters. For Andromeda and Centaurus A we fit an exponential disc and the DM halo, avoiding the central regions, with 5 free parameters. In these cases, the parameters are determined statistically, by applying a non-linear least-squares method, with initial parameters close to those reported in the literature.
@@
For the galaxies in Nashimoto et al. (2022), we found the rotation curves for 6 of the 9 galaxies observed. For those galaxies, we fit the rotation curves with mSFDM only. In this case, we have 3 free parameters only, $\rho_{0}$, $\rho_{1}$, and $l$. In this part, we use a Markov Chain Monte Carlo (MCMC) code implemented with the emcee library in Python. We report the errors at $1\sigma$.
III.1 Exponential Bulge Model
@@
For the bulge in the MW, we use the exponential-sphere model. In this case, the volume mass density $\rho_{\text{b}}$ is an exponential function of the radius $r$ with a scale radius $h_{b}$, as follows (Sofue, 2017):
$\rho_{\text{b}}(r)=\rho_{\text{b,0}}e^{-r/h_{b}},$ (35)
@@
with $\rho_{\text{b,0}}$ the central bulge density. The corresponding mass enclosed on a sphere of radius $r$ is
$M_{\text{b}}(r)=m_{\text{b}}\left[1-e^{-y}\left(1+y+\frac{y^{2}}{2}\right)\right],$ (36)
where $y:=r/h_{b}$, while the total mass of the bulge, $m_{\text{b}}$, is
$m_{\text{b}}=\int_{0}^{\infty}4\pi r^{2}\rho_{\text{b}}(r)\text{d}r=8\pi h_{b}^{3}\rho_{\text{b,0}}.$ (37)
@@
$v_{b}(r)=\sqrt{\frac{GM_{b}(r)}{r}}.$ (38)
III.2 Exponential disc model
For the galactic discs, we use the exponential Freeman model for the surface mass density (Freeman, 1970):
$\Sigma_{\text{d}}(r)=\Sigma_{\text{d,0}}e^{-r/h_{d}},$ (39)
where $\Sigma_{\text{d,0}}$ is the central density and $h_{d}$ is the scale radius. The total mass is given by $m_{\text{d}}=2\pi\Sigma_{\text{d,0}}h_{d}^{2}$. The corresponding rotation curve is given by (Binney and Tremaine, 2011)
$v_{\text{d}}(r)=\sqrt{4\pi G\Sigma_{\text{d,0}}h_{d}y^{2}\left[I_{0}(y)K_{0}(y)-I_{1}(y)K_{1}(y)\right]},$ (40)
where $y:=r/(2h_{d})$, and $I_{i}$ and $K_{i}$ are the modified Bessel functions.
IV Results
IV.1 Milky Way
@@
[CHECAR] The initial values for the least-squares method are taken from those reported in (Sofue, 2012): For the exponential bulge $m_{\text{b}}=(1.652\pm 0.083)\times 10^{10}M_{\odot}$ and $h_{\text{b}}=(522\pm 37)$ pc; for the exponential disc $m_{\text{d}}=(3.41\pm 0.41)\times 10^{10}M_{\odot}$ and $h_{d}=(3.19\pm 0.35)$ kpc. The total mass of the Milky Way, including the DM halo up to 150150~{}150150 kpc, has been estimated to $\sim 3\times 10^{11}M_{\odot}$ (Sofue, 2017). We did not take into account the gas mas contribution.
The best-fit parameters, obtained from a non-linear least-squares method implemented in Python for the exponential bulge and disc, and the ground and first excited states of the mSFDM halo, are listed in Table 1. In Figure 4, we show the best-fit MW rotation curve up to 70 kpc and the densities of the ground and first excited states of the mSFDM configuration. We cut the DM contribution as defined in (31).
Bulge $m_{b}=(1.163\pm 0.190)\times 10^{10}M_{\odot}$
$h_{b}=(176\pm 38)$ pc
Disc $m_{d}=(3.690\pm 1.250)\times 10^{10}M_{\odot}$
$h_{d}=(3.210\pm 1.035)$ kpc
mSFDM $l=(0.511\pm 0.072)\text{ kpc}^{-1}$
$\rho_{0}=(10.04\pm 4.27)\times 10^{-2}M_{\odot}\text{pc}^{-3}$
$\rho_{1}=(25.79\pm 12.63)\times 10^{-2}M_{\odot}\text{pc}^{-3}$
Table 1: Best-fit results for the Milky Way, from the least-squares method.
Figure 4: (Top) Best fit to the Milky Way with bulge, disc, and mSFDM components. Note that using the ground state and the first excited state is enough to fit the rotation curves of these galaxies very well. (Center) The same fit but now shows the central region of the MW galaxy. (Bottom) The behavior of the densities, using the same best-fit parameters.
IV.2 Andromeda
@@
For Andromeda, we exclude the central region from the analysis, because of the large surface density of baryons and a possible adiabatic contraction there (see Corbelli et al., 2010, and references therein). As the bulge is the most important contribution in that region, we do not fit such component, we use only an exponential disc, starting from 8.5 kpc. We use the data reported in that work.
@@
Following Corbelli et al. (2010), the disc luminosity used to obtain the mass parameter $m_{\text{d}}$ is $2.1\times 10^{10}L_{\odot}$. In their analysis, the authors vary the disc mass-to-light ratios between 2.5 and 8 $M_{\odot}/L_{\odot}$. For the best fitting $\Lambda$CDM model, they obtained $(M/L)_{d}=5.0M_{\odot}/L_{\odot}$, concentration parameter $C=12.0$, length-scale $h_{d}=4.5$ kpc (this parameter varies between 4.5 and 6.1 kpc), for an exponential disc. The total bulge + disc mass in that work is $1.3\times 10^{11}M_{\odot}$, and the dark NFW halo is $1.2\times 10^{12}M_{\odot}$.
@@
According to the last values, we use as initial numbers for our least-squares method: $m_{d}=(5.25-16.8)\times 10^{10}M_{\odot}$ and $h_{d}=4.5$ kpc.
For simplicity, we did not take into account the gas mass contribution.
The best-fit parameters for Andromeda are listed in Table 2. In Figure 5, we show the best-fit M31 rotation curve up to 36.5 kpc and the densities of the ground and first excited SF states.
Bulge $m_{b}=3.35\times 10^{10}M_{\odot}$
$h_{b}=430$ pc
Disc $m_{d}=5.01\times 10^{10}M_{\odot}$
$h_{d}=16.3$ kpc
mSFDM $l=(0.1534\pm 0.0054)\text{kpc}^{-1}$
$\rho_{0}=(3.100\pm 0.212)\times 10^{-2}M_{\odot}\text{pc}^{-3}$
$\rho_{1}=(0.140\pm 0.046)\times 10^{-2}M_{\odot}\text{pc}^{-3}$
Table 2: Best-fit results for M31, from the least-squares method.
Figure 5: (Top) Best-fit rotation curve of M31 with bulge, disc and mSFDM components. (Bottom) The behavior of the densities using the same best-fit parameters.
IV.3 Centaurus A

### v3

@@ 2.5 Exact Solution
See figure 2 for a 3D representation of the spherical ground state, the first excited state, and the sum of these two states of the SFDM. The main result of this work is that we obtained, in an analytical way, a spherical ground state, as expected, and a first excited state with a two-lobe shape towards the north and south of the center of the configuration. This structure is like that of the hydrogen atom, called a gravitational atom. The two-lobe structure can explain why satellite galaxies in MW-like galaxies are not distributed spherically, but rather along a north-south line, which has been called VPOS. This behavior has not only been observed in the Milky Way [6, 7], but also in Andromeda [8, 9] and Centaurus A [10, 11]. In this way, our result has the potential to explain this anisotropic distribution of satellite galaxies. Additionally, we tested the model with the distribution of satellites in 6 other galaxies reported in [65].
@@
Figure 2: (Left) A numerical simulation of the ground state of the SFDM. (Center) The same simulation but showing the evolution of the two-lobe excited state. (Right) The two states combined in the same image.
Now, from (2.29) and (2.30), the physical densities are given by $\rho_{k}(r)=R_{k}(r)R_{k}^{*}(r)$, and we define $\rho_{0}:=R_{00}^{2}$ and $\rho_{1}:=R_{10}^{2}$, with the constants $\rho_{0}$, the central density of the ground state, and $\rho_{1}$, the corresponding maximum density of the first excited state (see figure 3):
$\rho_{0}(r)=\rho_{0}\frac{\sin^{2}x}{x^{2}},$ (2.31)
$\rho_{1}(r)=\rho_{1}\left(\frac{\sin x}{x^{2}}-\frac{\cos x}{x}\right)^{2}.$ (2.32)
Notice that the density of the ground state (2.31) is of the same form as in the case of the real SF solution (equation (2.4)), for $j=1$ and $l=\pi/R$. In our case, the value $R=\pi/l$ defines the radius of the first minima of the ground state (see figure 3). In [43], the authors defined $R$ as the radius of the SFDM halo. However, in this work, since there is still contribution to the mass beyond the first minimum, as seen in figure 3, we define the maximum radius of every SF state, $R_{i}^{\text{max}}$, such that
$\rho_{i}(R_{i}^{\text{max}})=200\rho_{c},$ (2.33)
@@
Another difference with the previous work [43], is that the SFDM states in the case of the real SF are all centered at r=0, only their wavelength decreases and their amplitude changes, being all spherically symmetric. In that case, all the excited states share the definition of the radius $R$ of the halo. In the present work, for the complex SF, the second excited state has a different functional form, which gives it the two-lobe shape, with $\rho_{1}(0)=0$ and its maximum density outside $r=0$ (see figure 3).
@@
$M_{0}(r)=\frac{2\pi\rho_{0}}{l^{3}}\left(x-\cos x\sin x\right),$ (2.34)
$M_{1}(r)=\frac{2\pi\rho_{1}}{l^{3}}\left(\frac{\cos(2x)}{x}-\frac{1}{x}+\cos x\sin x+x\right).$ (2.35)
@@
Figure 3: The purple dashed line is the density of the ground state, centered at $r=0$. The blue solid line is the density of the first excited state, with its first maximum outside $r=0$. Note that the ground state might decay very quickly, and the first excited state decays slower, even if its first maximum is smaller than the first maximum of the ground state.
3 Rotation Curves of Galaxies
@@
To test this model and obtain the best values of the parameters $\rho_{0}$, $\rho_{1}$, and $l$ for the ground and the first excited states of the SFDM halo, respectively, we fit the rotation curves of the Milky Way, Andromeda, Centaurus A, as the main galaxies where the anisotropy of the satellite galaxies has been observed. Also, we fit the rotation curves of 6 MW-like galaxies whose satellite galaxies have been observed [65].
@@
The rotation curve (the tangential velocity of the satellites) is given by $v^{2}(r)=GM(r)/r$, for $M$ the mass of the galaxy. In what follows, we fit the rotation curves of different galaxies with this formula, such that
$v(r)=\sqrt{\frac{G}{r}\left(M_{b}+M_{d}+M_{0}+M_{1}\right)},$ (3.1)
@@
where the subscripts are for the bulge, disc, ground, and first excited states, respectively. We do not consider the gas contribution.
@@
To test the plausibility of the model in explaining the anisotropic distribution of satellite galaxies around a host galaxy, we fit the best models that reproduce the observed rotation curve of the main galaxy, with the ground state (spherically symmetric) and the first excited state (with two lobes) of the mSFDM as the DM halo. As said before, we define the radius of every state such that $\rho_{i}(R_{i})=200\rho_{c}$, for $i=0,1$, where $\rho_{c}=1.283\times 10^{-7}M_{\odot}\text{pc}^{-3}$ is the critical density of the Universe (equation (2.33)). Both states present density oscillations until reaching the density of $200\rho_{c}$. These halos are like concentric shells that contribute less and less to the gravitational potential.
@@
We assume that the motion of stars and gas in the galaxies is circular, without considering non-axisymmetric structures such as the bar and spiral arms. For the Milky Way, we fit an exponential bulge, an exponential disc, and the mSFDM halo, through 7 free parameters. For Andromeda, Centaurus A, and NGC 3437, we fit an exponential disc and the mSFDM halo, avoiding the central regions, through 5 free parameters, with initial parameters close to those reported in the literature.
For the last 5 galaxies from the work of [65] (NGC 2950, 3245, 3338, 5866, and 7332), we fit the corresponding rotation curves with mSFDM only. In these cases, we have 3 free parameters only, $\rho_{0}$, $\rho_{1}$, and $l$.
@@
To fit the free parameters as explained before, we implemented a Markov Chain Monte Carlo (MCMC) code with the emcee library in Python [66], as explained in Section 4. We report the best-fit parameters and errors in Section 6.
3.1 Exponential Bulge Model
@@
For the bulge in the MW, we use the exponential-sphere model. In this case, the volume mass density $\rho_{b}$ is an exponential function of the radius $r$ with a scale radius $h_{b}$, as follows [67]:
$\rho_{b}(r)=\rho_{b,0}e^{-r/h_{b}},$ (3.2)
@@
with $\rho_{b,0}$ the central bulge density. The corresponding mass enclosed on a sphere of radius $r$ is
$M_{b}(r)=m_{b}\left[1-e^{-y}\left(1+y+\frac{y^{2}}{2}\right)\right],$ (3.3)
where $y:=r/h_{b}$, while the total mass of the bulge, $m_{b}$, is
$m_{b}=\int_{0}^{\infty}4\pi r^{2}\rho_{b}(r)\text{d}r=8\pi h_{b}^{3}\rho_{b,0}.$ (3.4)
@@
$v_{b}(r)=\sqrt{\frac{GM_{b}(r)}{r}}.$ (3.5)
3.2 Exponential Disc Model
For the galactic discs, we use the exponential Freeman model for the surface mass density [68]:
$\Sigma_{d}(r)=\Sigma_{d,0}e^{-r/h_{d}},$ (3.6)
where $\Sigma_{d,0}$ is the central density and $h_{d}$ is the scale radius. The total mass is given by $m_{d}=2\pi\Sigma_{d,0}h_{d}^{2}$. The corresponding rotation curve is given by
$v_{d}(r)=\sqrt{4\pi Gy^{2}\Sigma_{d,0}h_{d}\left[I_{0}(y)K_{0}(y)-I_{1}(y)K_{1}(y)\right]},$ (3.7)
where $y:=r/(2h_{d})$, and $I_{i}$ and $K_{i}$ (for $i=0,1$) are the modified Bessel functions.
4 MCMC Calibration Method
We use a Markov Chain Monte Carlo (MCMC) code to constrain the free parameters of the mSFDM model and find the best fit for the velocity rotation curve in each galaxy. We implemented the affine invariant ensemble sampler [69] for MCMC using the emcee library in Python [66], maximizing the likelihood function $\mathcal{L}(\mathbf{p})$ given by
$\mathcal{L}({\bf p})=\frac{1}{(2\pi)^{N/2}|{\bf C}|^{1/2}}\exp{\left(-\frac{{\bf\Delta}^{T}{\bf C}^{-1}{\bf\Delta}}{2}\right)},$ (4.1)
where $\mathbf{p}$ is the vector of parameters, $N$ the number of data points for each galaxy, ${\bf\Delta}=v_{\mathrm{obs}}(r_{i})-v(r_{i},\mathbf{p})$, for $v_{\mathrm{obs}}$ the observational circular velocity at radius $r_{i}$, and $v(r_{i},\mathbf{p})$ the total velocity given by (3.1), computed in the same position where $v_{\mathrm{obs}}$ was measured, and $\mathbf{C}$ a diagonal matrix.
We sample the parameter space from uniform prior ranges using $50$ random walkers, each with $5\times 10^{4}$ steps. We tested the convergence of the fit by monitoring the mean integrated autocorrelation time $\tau_{\text{int}}$ at each step. Convergence is considered achieved when the number of steps $N_{\text{step}}$ is at least 100 times $\tau_{\text{int}}$, and $\tau_{\text{int}}$ reaches an approximately steady state, fluctuating by less than $1\%$ per step (see references [70, 69, 66], for a complete discussion of the use of autocorrelation time as a convergence criterion).
We estimate the best-fit parameters by calculating the median of the posterior distribution associated with each parameter. We present the $1D$ histograms and $2D$ contours of the posterior distributions, as well as the value $-2\ln\mathcal{L}$ associated with the best fit. The fitting parameters and $1\sigma$ and $2\sigma$ confidence levels (CL) are computed from the Markov chains with 30% as burn-in. We plotted the histograms and contour plots using the Python package GetDist[71].
5 Observational Data for the MCMC method
In table 1, we present a summary of the characteristics of the galaxies used in the present work, as well as the references from which we obtained the data for this article. In the following subsections, we give more details of the host galaxies studied.
In Appendix A, in Table 4, we report the priors used for the MCMC calibration method, for each one of the galaxies and the mSFDM and halo components.
Host Galaxy Morphology i $v_{\text{max}}$ $D$ $r_{\text{max}}$ Source
(°) (km/s) (Mpc) (kpc)
MW SB 250 0.008* 64.5 [67]
M31 Sb* 72.2* 254 0.75* 36.5 [72]
Cen A S0* 45.3* 257 4.0* 6.3 [73]
NGC 2950 SB0 50 156 18.9 1.01 [74]
NGC 3245 S0* 61.8* 147 20.9* 1.34 [75]
NGC 3338 Sc 55 174 23.1* 40.7 [76]
NGC 3437 SABc 69 194 22.3* 4.47 [77]
NGC 5690 Sc* 75.9* 183 19.1* 7.55 [77]
NGC 5866 SA0 90* 210 14.7* 3.4 [78]
NGC 7332 S0 90* 147 23.0* 5.3 [79]
Table 1: Characteristics of the galaxies used in the present work. The columns read: (1) The name of the galaxy, (2) Morphology, (3) Inclination between the line-of-sight and the polar axis of the galaxy, (4) Maximum velocity of the rotation curve, $v_{\text{max}}$, (5) Distance to the galaxy, $D$, (6) Maximum spatial extent in kpc of either the approaching or receding side of the galaxy, $r_{\text{max}}$, (7) Source of the data used in this work (for the data with * the source is HyperLeda database [80, http://leda.univ-lyon1.fr/] and references therein).
5.1 Milky Way
@@
The initial values for the MCMC method are taken from those reported in [82]: For the exponential bulge $m_{b}=(1.652\pm 0.083)\times 10^{10}M_{\odot}$ and $h_{b}=(522\pm 37)$ pc; for the exponential disc $m_{d}=(3.41\pm 0.41)\times 10^{10}M_{\odot}$ and $h_{d}=(3.19\pm 0.35)$ kpc. The total mass of the Milky Way, including the DM halo up to $\sim 150$ kpc, has been estimated to be $\sim 3\times 10^{11}M_{\odot}$ [67]. We did not take into account the gas contribution.
5.2 Andromeda
@@
For Andromeda, we exclude the central region from the analysis, because of the large surface density of baryons and a possible adiabatic contraction there [see 72, and references therein]. As the bulge is the most important contribution in that region, we do not fit such a component, we use only an exponential disc, starting from 8.5 kpc.
@@
We use the data reported in [72]. Following that work, the disc luminosity $L_{d}$ used to obtain the mass parameter $m_{d}$ is $2.1\times 10^{10}L_{\odot}$. In their analysis, the authors vary the disc mass-to-light ratios between 2.5 and 8 $M_{\odot}/L_{\odot}$. For the best fitting $\Lambda$CDM model, they obtained $(M/L)_{d}=5.0M_{\odot}/L_{\odot}$, concentration parameter $C=12.0$, and length-scale $h_{d}=4.5$ kpc (this parameter varies between 4.5 and 6.1 kpc), for an exponential disc. The total bulge + disc mass in that work is $1.3\times 10^{11}M_{\odot}$, and the dark NFW halo is $1.2\times 10^{12}M_{\odot}$.
@@
According to the last values, we use as initial numbers for the MCMC method: $m_{d}=(5.25-16.8)\times 10^{10}M_{\odot}$ and $h_{d}=4.5$ kpc. For simplicity, we did not take into account the gas mass contribution.
5.3 Centaurus A